# When and Where do Data Poisons Attack Textual Inversion?

### v2

@@ Abstract
**footnotetext: Equal contribution
@@ 1 Introduction
Unfortunately, the surge in generative AI corresponds to misfortune for artists and copyright owners, who may see their works easily reproduced by retrained generative models. Consequently, the subfield of data poisoning for generative AI has focused on how to inject images with signals that nullify their use in training datasets for generative models. Individuals looking to protect their data are encouraged to “poison" their images with imperceptible adversarial perturbations such that generative models will fail to learn to recreate the novel concepts contained within.
Figure 1: Applying SZT to TI mitigates poison signals and recovers shape, color, and texture of desired concepts. NovelConcepts10 training data poisoned by ADM+. Generation prompt: “A $R*$ with a wheat field in the background"
@@ 1.1 Research Gap and Contributions
In this paper, we examine the learning behavior of TI [15], a lightweight personalization technique for adapting DMs to novel concepts, when trained on poisoned datasets. To enable our analysis, we devise a novel visualization method, semantic sensitivity maps (SSM), to display loss sensitivity to individual text tokens throughout training. Using SSMs, we observe several heretofore unseen trends in poison learning: (1) DMs concentrate poison learning at lower-middle timesteps (approximately $t\in[0,600]$) and (2) adversarial signals tend to “distract" learning away from the novel concept of interest and towards irrelevant regions.
@@ 2 Related Works
Although DMs can easily generate most concepts from their training dataset with high quality, they predictably suffer in generation quality for novel (unseen) concepts. Editing techniques [50, 20, 52, 51, 29, 82, 3] apply modifications to an input image while enforcing structural or composition constraints, permitting generation with novel concept images. However, editing techniques are still bound to the prior from the DM’s training data. In order to “inject" new concepts and styles into the DM, personalization techniques [25, 15, 19, 83, 74, 53, 59, 35, 73, 18] quickly finetune some component of the DM on a small dataset of novel concept images. Many personalization techniques learn new text embeddings [15, 18, 35, 73, 74] or co-opt existing tokens [59, 19] to capture concept semantics. Without retraining the entire DM from scratch, personalization techniques are the best way to adapt DMs to new concepts. In particular, TI [15] is a lightweight personalization method that requires training only one new text embedding and is often included within other methods [35, 73, 18].
@@
AEs for DMs fall into two main categories. Encoder attacks [63, 64, 61] utilize the encoder of an LDM to learn a small perturbation to an input image that minimizes the distance between the perturbed image encoding and the encoding of some unrelated target image. Diffuser attacks [37, 39, 78, 61, 6] aim to learn a small perturbation to an input image that maximizes the DM loss (often noise-prediction). Other attack methods combine encoder attacks and diffuser attacks [38, 61]. Defenses against AEs for DMs include augmentation [65] and regeneration [13, 77, 84, 44], where poisoned images are randomly noised and then denoised with a DM to remove adversarial signals. Regeneration methods are the most actively researched, but they are time-consuming and complex as they require preprocessing entire datasets with a DM in order to “purify" them.
@@ 3.1 Latent Diffusion Models
where $\overline{\alpha}_{t}=\Sigma_{s=1}^{t}(1-\beta_{s})$ and $\epsilon\sim\mathcal{N}(\textbf{0},I)$. $\overline{\alpha}_{t}$ monotonically decreases from $1$ to 00 as $t$ increases from 00 to $T$, following a variance-preserving schedule [68]. A generative DM learns the denoising process, where the goal is to maximize the likelihood of generating the unnoised input ($z_{0}$). DDPM [24] showed that this objective can be practically realized with a simple noise-prediction task. The model predicts the Gaussian noise $\epsilon$ added to input $z_{0}$ when given noisy sample $z_{t}$, and the learning objective is given as
$\operatorname*{arg\,min}_{\theta}\ E_{x,t\sim\mathcal{U}(0,T),\epsilon}\left[||\epsilon_{\theta}(z_{t},t)-\epsilon||_{2}^{2}\right],$ (2)
where $x$ is drawn from the training dataset, $\mathcal{U}(0,T)$ is the uniform distribution between timesteps 00 and $T$, and $\epsilon_{\theta}$ is a model that predicts a noise vector $\epsilon$ given a noised latent $z_{t}$. For text-conditional models, $\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))$ also accepts features vector $\tau_{\theta}(y)$, where $\tau_{\theta}$ is a language model and input prompt $y$ is provided by the user or training data.
@@ 3.2 Textual Inversion
$\operatorname*{arg\,min}_{\theta_{R^{*}}}E_{x_{f},t\sim\mathcal{U}(0,T),y^{*},\epsilon}\left[||\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y^{*}))-\epsilon||_{2}^{2}\right],$ (3)
@@ 3.3 Adversarial Examples
$\begin{split}\operatorname*{arg\,max}_{\delta}&\ E_{x_{p},t\sim\mathcal{U}(0,T),\epsilon}\left[||\epsilon_{\theta}(z^{\prime}_{t},t)-\epsilon||_{2}^{2}\right]\\ &\text{s.t.}\quad||\delta||_{b}\leq\kappa,\end{split}$ (4)
@@
$\operatorname*{arg\,min}_{\delta}\ E_{x_{p}}\left[||\mathcal{E}(x_{p}+\delta)-\mathcal{E}(x_{g})||_{2}^{2}\right]\ \text{s.t.}\ ||\delta||_{b}\leq\kappa.$ (5)
@@ 4.1 Visualizing the Effect of Adversarial Examples
To observe the influence of AEs against TI, it is critical to visualize the impact of each text embedding on a generated image. Let $\widehat{e}=\{e_{0},...e_{n},...,e_{L-1}\}$, where $e_{0}$ represents the start token embedding vector and $e_{n}$ represents the embedding vector of the $n$th token. For a text input consisting of $l$ tokens, $\widehat{e}$ is padded to length $L$ with $\{e_{l+1},...,e_{L-1}\}$ with end token embeddings. Each $e_{n}$ is obtained via lookup in a token-embedding dictionary. As an example, given the text prompt, “a puppy wearing a hat", embeddings $\{e_{1},...,e_{5}\}$ correspond to “a", “puppy", “wearing", “a", and “hat" respectively.
A popular text-attribution method for LDMs visualizes the cross-attention maps [41] in $\epsilon_{\theta}$. However, cross-attention maps are incapable of capturing the effect of each embedding $e_{n}$. This is because the self-attention layers in the text encoder $\tau_{\theta}$ entangle signals from the embedding vectors of $\widehat{e}$ before they reach the cross-attention layers of $\epsilon_{\theta}$. Fig. 2(a) visualizes text attribution using cross-attention maps. The cross-attention map corresponding to the start token $e_{0}$ is bright, highlighting many regions, but the maps corresponding to words “puppy" and “hat" are dim and blurry. This drawback has been noted in other works [46].
Figure 2: Cross attention maps (a) and SSMs (b) for tokens “<start _token>", “puppy", and “hat" with text prompt “a puppy wearing a hat" at noise timestep $t=500$.
@@
$SSM(x,t,\widehat{e},n)$ $=E_{\widehat{e}_{\Delta n}}\left[||\epsilon_{\theta}(z_{t},t,c)-\epsilon_{\theta}(z_{t},t,c_{\Delta n})||_{ch}^{2}\right]$
@@
where $x$ is an input image and $c$ is the condition tensor. $c_{\Delta n}$ is analogous to $c$, except that it is derived from $\widehat{e}_{\Delta n}$, a modified version of $\widehat{e}$ that replaces $e_{n}$ with a randomly sampled vector from the embedding dictionary. For notational simplicity, we denote that $\tau_{\theta}(\widehat{e})$ accepts token embeddings as input instead of the prompt. $||\cdot||_{ch}^{2}$ indicates sum of squares across channels such that the output is a one-channel image that matches the input height and width. Fig. 2(b) displays the SSMs corresponding to the embeddings for “puppy" and “hat". Compared to cross-attention attribution maps in Fig. 2(a), SSMs more accurately reflect the regions corresponding to each input embedding. Crucially, SSMs avoid attributing semantic tokens to the start token.
Using SSMs, we can investigate when and where adversarial signals attack the TI process. Fig. 3 shows the SSMs for a clean image and its corresponding AEs for various noising timesteps. The input images contain a novel concept not contained in the training dataset. The image prompt is “a $R^{*}$ laying on top of a grass covered field". Fig. 3 shows the SSMs for the trainable text embedding corresponding to $R^{*}$ before and after TI training. We observe that before TI training, SSMs of the clean sample and AEs highlight the novel object in the range $t\in[300,700]$. This indicates that the DM prior can already correlate unknown objects with unknown tokens. After TI training, the novel concept is more pronounced in SSMs of clean images, but the SSMs of AEs become extremely noisy, particularly at lower-middle timestep values. Notably, the attribution for AEs extends far outside the novel concept region. To explain this behavior, we investigate the temporal and spatial learning properties of TI in subsequent sections.
@@ 4.2 Timestep Learning Bias
Figure 4: Loss distributions for Stable Diffusion 1.5 on $1000$ clean (green) and $1000$ ADM+ poisoned (red) LAION Aesthetic images at each of $21$ interspersed timesteps. The dark lines represent median loss values. The dark and light bands encompass 50% and 95%, respectively, of the loss distributions.
@@
Recalling that $\overline{\alpha}_{t}$ from Eq. 1 monotonically decreases from $1$ at $t=0$ to 00 at $t=T$, we can see that the noising process is interpolating between input $z_{0}$ and noise $\epsilon$. At $t=0$, $z_{t=0}=\sqrt{\overline{\alpha}_{0}}z_{0}+\sqrt{1-\overline{\alpha}_{0}}\epsilon=z_{0}$. Therefore, at $t=0$, the input to the noise-prediction model $\epsilon_{\theta}(z_{0},0)$ contains no information about the sampled noise $\epsilon$. Since $\epsilon$ is distributed as $\mathcal{N}(\textbf{0},I)$, the optimal noise prediction to minimize the expected loss at $t=0$ is $\epsilon_{\theta}(z_{0},0)=\textbf{0}$. Predictably, the loss at $t=0$ is a maximum because the noise-prediction model cannot accurately predict the ground-truth noise. Conversely, at $t=T$, $z_{t=T}=\sqrt{\overline{\alpha}_{T}}z_{0}+\sqrt{1-\overline{\alpha}_{T}}\epsilon=\epsilon$, and the noise-prediction model simply needs to return the input to minimize loss. At $t=T$, the loss approaches a minimum.
@@
To validate our argument, we calculate the noise-prediction loss of Stable Diffusion 1.5 for $1000$ images sampled from the LAION Aesthetic [62] dataset at $21$ equally-spaced timesteps ($21000$ samples total) and plot the resulting distributions in Fig. 4. As predicted, the expected loss decreases from a maximum at $t=0$ to 00 at $t=T$. Applying ADM+ [39] poisoning to the images increases the loss values in the lower-middle timesteps (as expected by a loss-maximizing AE algorithm), but the losses at $t=0$ and at high timesteps remain unchanged.
@@
We must also consider how the noise-prediction objective affects gradients during TI training. Fig. 5 displays the average loss gradient magnitude with respect to the trainable text embedding at each timestep after $1000$ steps of TI training with Stable Diffusion 1.5. Curves are obtained by calculating $1000$ losses for each concept of the NovelConcepts10 dataset across $21$ equally-spaced timesteps and then averaging gradient magnitudes at each timestep across concepts. As expected, the graph shows that at high timesteps, gradients decay to 00 as the loss distributions converge to 00. At lower-middle timesteps, moderate loss allows the model to backpropagate real (or adversarial) signals to the text embedding, corresponding to high gradients. However, near $t=0$, high losses correspond to low gradients. Intuitively, this is because the noise-prediction model minimizes loss at $t=0$ by predicting near-0 noise regardless of the inputs, $z_{t}$ and $\tau_{\theta}(y^{*})$. Since the prediction is independent of the inputs, no loss signal is backpropagated to the trainable text embedding. The loss gradient magnitudes in Fig. 5 correspond to the intensities of the sensitivity maps in Fig. 3.
@@ 4.3 Spatial Learning Bias
As observed in Sec. 4.1, AEs disrupt TI by “distracting" the optimization process and expanding the spatial distribution of the trainable token attribution beyond the novel concept region. To understand this phenomenon, we revisit the adversarial objectives in Eq. 4 and Eq. 5.
We begin by noting that both adversarial objectives involve an $l^{2}$ norm which includes a sum over all spatial coordinates (height, width, channels) of the error tensor. This summation does not apply preferential weight to any spatial region. Furthermore, the $l^{2}$ vector norm is convex and non-decreasing, implying that it increases as the magnitudes of individual elements increase. This encourages all spatial elements to contribute to optimizing the $l^{2}$ objective. This incentive combines with the adversarial perturbation constraint $||\delta||_{b}$, which is commonly implemented with the $l^{\infty}$ norm. We note that $||\delta||_{\infty}$ is non-increasing for all infinitesimal changes in perturbation elements $\delta_{i}$ where $|\delta_{i}|<||\delta||_{\infty}$; this permits increases in most perturbation elements without affecting the constraint. In summary, the $l^{2}$ norm objectives in Eq. 4 and Eq. 5 encourage all elements of the error terms to contribute and the $l^{\infty}$ constraint on $\delta$ encourages all elements of $\delta$ to contribute maximally. Practically, this implies that adversarial perturbations maximally infect every region of a poisoned image.
@@
Figure 6: Loss difference inside (“LossIn") and outside (“LossOut") novel concept regions during TI training on clean, ADM+, SDS+, and EA images of seven concepts from NovelConcepts10.
@@ 4.5 Safe-Zone Training
$\operatorname*{arg\,min}_{\theta_{R^{*}}}E_{x^{\prime}_{j},t,y^{*},\epsilon}\left[||(\epsilon_{\theta}(z^{\prime}_{t},t,\tau_{\theta}(y^{*}))-\epsilon)\odot M_{z}||_{2}^{2}\right]$ (7)
@@ 5.2.2 Timestep Range Ablation
Based on findings about timestep learning bias in Sec. 4.2, we ablate methods to restrict timesteps to high values during TI training. We consider multiple sampling methods to favor high timesteps: thresholding, power distributions, and $tanh$ distributions. Fig. 15 in Appendix H displays the probability curves for the power and $tanh$ distributions. We ablate $t$-sampling methods for clean, ADM+, ADM-, and EA versions of NovelConcepts10. Full results are given in Appendix H. We find that simple high-thresholding works best, improving DINOv2 similarity for ADM+ data from $0.10$ ($t\sim\mathcal{U}(0,1000))$ to $0.21$ for $t\geq 600$ and $0.28$ for $t\geq 700$. Using $t\geq 600$ also maintains or improves scores for clean and EA datasets. In our subsequent proposed defenses, we mainly utilize $t\geq 600$, denoted “T600".

### v4

@@ Abstract
**footnotetext: Equal contribution, may cite either first
@@ 1 Introduction
Unfortunately, the surge in generative AI corresponds to misfortune for artists and copyright owners, who may see their works easily reproduced by retrained generative models. Consequently, the subfield of data poisoning for generative AI has focused on how to inject images with signals that nullify their use in training datasets for generative models. Individuals looking to protect their data are encouraged to “poison” their images with imperceptible adversarial perturbations such that generative models will fail to learn to recreate the novel concepts contained within.
Figure 1: Applying SZT to TI mitigates poison signals and recovers shape, color, and texture of desired concepts. NovelConcepts10 training data poisoned by ADM+. Generation prompt: “A RR* with a wheat field in the background”
@@ 1.1 Research Gap and Contributions
In this paper, we examine the learning behavior of TI [15], a lightweight personalization technique for adapting DMs to novel concepts, when trained on poisoned datasets. To enable our analysis, we devise a novel visualization method, semantic sensitivity maps (SSM), to display loss sensitivity to individual text tokens throughout training. Using SSMs, we observe several heretofore unseen trends in poison learning: (1) DMs concentrate poison learning at lower-middle timesteps (approximately $t\in[0,600]$) and (2) adversarial signals tend to “distract” learning away from the novel concept of interest and towards irrelevant regions.
@@ 2 Related Works
Although DMs can easily generate most concepts from their training dataset with high quality, they predictably suffer in generation quality for novel (unseen) concepts. Editing techniques [50, 20, 52, 51, 29, 82, 3] apply modifications to an input image while enforcing structural or composition constraints, permitting generation with novel concept images. However, editing techniques are still bound to the prior from the DM’s training data. In order to “inject” new concepts and styles into the DM, personalization techniques [25, 15, 19, 83, 74, 53, 59, 35, 73, 18] quickly finetune some component of the DM on a small dataset of novel concept images. Many personalization techniques learn new text embeddings [15, 18, 35, 73, 74] or co-opt existing tokens [59, 19] to capture concept semantics. Without retraining the entire DM from scratch, personalization techniques are the best way to adapt DMs to new concepts. In particular, TI [15] is a lightweight personalization method that requires training only one new text embedding and is often included within other methods [35, 73, 18].
@@
AEs for DMs fall into two main categories. Encoder attacks [63, 64, 61] utilize the encoder of an LDM to learn a small perturbation to an input image that minimizes the distance between the perturbed image encoding and the encoding of some unrelated target image. Diffuser attacks [37, 39, 78, 61, 6] aim to learn a small perturbation to an input image that maximizes the DM loss (often noise-prediction). Other attack methods combine encoder attacks and diffuser attacks [38, 61]. Defenses against AEs for DMs include augmentation [65] and regeneration [13, 77, 84, 44], where poisoned images are randomly noised and then denoised with a DM to remove adversarial signals. Regeneration methods are the most actively researched, but they are time-consuming and complex as they require preprocessing entire datasets with a DM in order to “purify” them.
@@ 3.1 Latent Diffusion Models
where $\overline{\alpha}_{t}=\Sigma_{s=1}^{t}(1-\beta_{s})$ and $\epsilon\sim\mathcal{N}(\textbf{0},I)$. $\overline{\alpha}_{t}$ monotonically decreases from $1$ to 0 as $t$ increases from 0 to $T$, following a variance-preserving schedule [68]. A generative DM learns the denoising process, where the goal is to maximize the likelihood of generating the unnoised input ($z_{0}$). DDPM [24] showed that this objective can be practically realized with a simple noise-prediction task. The model predicts the Gaussian noise $\epsilon$ added to input $z_{0}$ when given noisy sample $z_{t}$, and the learning objective is given as
[…]
where $x$ is drawn from the training dataset, $\mathcal{U}(0,T)$ is the uniform distribution between timesteps 0 and $T$, and $\epsilon_{\theta}$ is a model that predicts a noise vector $\epsilon$ given a noised latent $z_{t}$. For text-conditional models, $\epsilon_{\theta}(z_{t},t,\tau_{\theta}(y))$ also accepts features vector $\tau_{\theta}(y)$, where $\tau_{\theta}$ is a language model and input prompt $y$ is provided by the user or training data.
@@ 3.2 Textual Inversion
[…]
@@ 3.3 Adversarial Examples
[…]
@@
[…]
@@ 4.1 Visualizing the Effect of Adversarial Examples
To observe the influence of AEs against TI, it is critical to visualize the impact of each text embedding on a generated image. Let $\widehat{e}=\{e_{0},...e_{n},...,e_{L-1}\}$, where $e_{0}$ represents the start token embedding vector and $e_{n}$ represents the embedding vector of the $n$th token. For a text input consisting of $l$ tokens, $\widehat{e}$ is padded to length $L$ with $\{e_{l+1},...,e_{L-1}\}$ with end token embeddings. Each $e_{n}$ is obtained via lookup in a token-embedding dictionary. As an example, given the text prompt, “a puppy wearing a hat”, embeddings $\{e_{1},...,e_{5}\}$ correspond to “a”, “puppy”, “wearing”, “a”, and “hat” respectively.
A popular text-attribution method for LDMs visualizes the cross-attention maps [41] in $\epsilon_{\theta}$. However, cross-attention maps are incapable of capturing the effect of each embedding $e_{n}$. This is because the self-attention layers in the text encoder $\tau_{\theta}$ entangle signals from the embedding vectors of $\widehat{e}$ before they reach the cross-attention layers of $\epsilon_{\theta}$. Fig. 2(a) visualizes text attribution using cross-attention maps. The cross-attention map corresponding to the start token $e_{0}$ is bright, highlighting many regions, but the maps corresponding to words “puppy” and “hat” are dim and blurry. This drawback has been noted in other works [46].
Figure 2: Cross attention maps (a) and SSMs (b) for tokens “<start _token>”, “puppy”, and “hat” with text prompt “a puppy wearing a hat” at noise timestep $t=500$.
@@
$\text{s.t.}\ c$ $=\tau_{\theta}(\widehat{e}),\ c_{\Delta n}=\tau_{\theta}(\widehat{e}_{\Delta n}),$ (6)
[…]
Using SSMs, we can investigate when and where adversarial signals attack the TI process. Fig. 3 shows the SSMs for a clean image and its corresponding AEs for various noising timesteps. The input images contain a novel concept not contained in the training dataset. The image prompt is “a $R^{*}$ laying on top of a grass covered field”. Fig. 3 shows the SSMs for the trainable text embedding corresponding to $R^{*}$ before and after TI training. We observe that before TI training, SSMs of the clean sample and AEs highlight the novel object in the range $t\in[300,700]$. This indicates that the DM prior can already correlate unknown objects with unknown tokens. After TI training, the novel concept is more pronounced in SSMs of clean images, but the SSMs of AEs become extremely noisy, particularly at lower-middle timestep values. Notably, the attribution for AEs extends far outside the novel concept region. To explain this behavior, we investigate the temporal and spatial learning properties of TI in subsequent sections.
@@ 4.2 Timestep Learning Bias
Figure 4: Loss distributions for Stable Diffusion 1.5 on $1000$ clean (green) and $1000$ ADM+ poisoned (red) LAION Aesthetic images at each of 2121 interspersed timesteps. The dark lines represent median loss values. The dark and light bands encompass 50% and 95%, respectively, of the loss distributions.
@@
Recalling that $\overline{\alpha}_{t}$ from Eq. 1 monotonically decreases from $1$ at $t=0$ to 0 at $t=T$, we can see that the noising process is interpolating between input $z_{0}$ and noise $\epsilon$. At $t=0$, $z_{t=0}=\sqrt{\overline{\alpha}_{0}}z_{0}+\sqrt{1-\overline{\alpha}_{0}}\epsilon=z_{0}$. Therefore, at $t=0$, the input to the noise-prediction model $\epsilon_{\theta}(z_{0},0)$ contains no information about the sampled noise $\epsilon$. Since $\epsilon$ is distributed as $\mathcal{N}(\textbf{0},I)$, the optimal noise prediction to minimize the expected loss at $t=0$ is $\epsilon_{\theta}(z_{0},0)=\textbf{0}$. Predictably, the loss at $t=0$ is a maximum because the noise-prediction model cannot accurately predict the ground-truth noise. Conversely, at $t=T$, $z_{t=T}=\sqrt{\overline{\alpha}_{T}}z_{0}+\sqrt{1-\overline{\alpha}_{T}}\epsilon=\epsilon$, and the noise-prediction model simply needs to return the input to minimize loss. At $t=T$, the loss approaches a minimum.
@@
To validate our argument, we calculate the noise-prediction loss of Stable Diffusion 1.5 for $1000$ images sampled from the LAION Aesthetic [62] dataset at $21$ equally-spaced timesteps ($21000$ samples total) and plot the resulting distributions in Fig. 4. As predicted, the expected loss decreases from a maximum at $t=0$ to 0 at $t=T$. Applying ADM+ [39] poisoning to the images increases the loss values in the lower-middle timesteps (as expected by a loss-maximizing AE algorithm), but the losses at $t=0$ and at high timesteps remain unchanged.
@@
We must also consider how the noise-prediction objective affects gradients during TI training. Fig. 5 displays the average loss gradient magnitude with respect to the trainable text embedding at each timestep after $1000$ steps of TI training with Stable Diffusion 1.5. Curves are obtained by calculating $1000$ losses for each concept of the NovelConcepts10 dataset across $21$ equally-spaced timesteps and then averaging gradient magnitudes at each timestep across concepts. As expected, the graph shows that at high timesteps, gradients decay to 0 as the loss distributions converge to 0. At lower-middle timesteps, moderate loss allows the model to backpropagate real (or adversarial) signals to the text embedding, corresponding to high gradients. However, near $t=0$, high losses correspond to low gradients. Intuitively, this is because the noise-prediction model minimizes loss at $t=0$ by predicting near-0 noise regardless of the inputs, $z_{t}$ and $\tau_{\theta}(y^{*})$. Since the prediction is independent of the inputs, no loss signal is backpropagated to the trainable text embedding. The loss gradient magnitudes in Fig. 5 correspond to the intensities of the sensitivity maps in Fig. 3.
@@ 4.3 Spatial Learning Bias
As observed in Sec. 4.1, AEs disrupt TI by “distracting” the optimization process and expanding the spatial distribution of the trainable token attribution beyond the novel concept region. To understand this phenomenon, we revisit the adversarial objectives in Eq. 4 and Eq. 5.
[…]
@@
Figure 6: Loss inside (“LossIn”) and outside (“LossOut”) novel concept regions during TI training on clean, ADM+, SDS+, and EA data, averaged across all concepts in NovelConcepts10.
@@ 4.5 Safe-Zone Training
$\text{s.t.}\ x^{\prime}_{j}=JPEG(x^{\prime}_{p}),\ t\sim\mathcal{U}(t_{th},T),\ M_{z}=Resize(M_{x}),$
@@ 5.2.2 Timestep Range Ablation
Based on findings about timestep learning bias in Sec. 4.2, we ablate methods to restrict timesteps to high values during TI training. We consider multiple sampling methods to favor high timesteps: thresholding, power distributions, and tanhtanh distributions. Fig. 15 in Appendix H displays the probability curves for the power and tanhtanh distributions. We ablate $t$-sampling methods for clean, ADM+, ADM-, and EA versions of NovelConcepts10. Full results are given in Appendix H. We find that simple high-thresholding works best, improving DINOv2 similarity for ADM+ data from $0.10$ ($t\sim\mathcal{U}(0,1000))$ to $0.21$ for $t\geq 600$ and $0.28$ for $t\geq 700$. Using $t\geq 600$ also maintains or improves scores for clean and EA datasets. In our subsequent proposed defenses, we mainly utilize $t\geq 600$, denoted “T600”.